Let X_1, X_2, ..., X_21 be a random sample from a normal distribution with µ = 10, σ^2 = 4. Using the chi-square distribution, find P(2 < S^2 < 8)
Answer #1

for we know that (n-1)s2/ σ 2 follows chi distribution with n-1 degree of freedom

therefore P(2<S2<8)=P(2*(21-1)/4<(n-1)*s22 <8*(21-1)/4)=P(10<X2 <40)=0.9950-0.0318=0.9632